OpenStudy (anonymous):

Find the derivative using implicit differentiation ln(7+e^xy)=y

hartnn (hartnn):

before differentiating, you can simplify it, use, if \(\large \ln (a)=b \implies a=e^b\)

OpenStudy (anonymous):

What would that be

hartnn (hartnn):

then, \(\large \ln(7+e^{xy})=y \implies 7+e^{xy}=e^y\) right ? now you can take the derivative.

OpenStudy (anonymous):

but I keep getting the wrong answer I got ye=7

hartnn (hartnn):

whats ye ?

OpenStudy (anonymous):

y^e-7

hartnn (hartnn):

but you'll need product rule (also, chain rule) for left side, could you observe that ? for the term e^{xy}